OpenStudy (anonymous):

a+b+c=16, what is b?

OpenStudy (anonymous):

well whats c and a?

OpenStudy (anonymous):

is this multiple choice

OpenStudy (anonymous):

im gonna try to upload a pic of the problem

OpenStudy (xapproachesinfinity):

you want to solve for b?

OpenStudy (xapproachesinfinity):

in terms of a and c

OpenStudy (anonymous):

@xapproachesinfinity and @KlOwNlOvE i uploaded a pic of the problem

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @thomaster @goformit100 @vzfreakz can one of yall help

OpenStudy (anonymous):

@KlOwNlOvE you now understand the problem, right?

OpenStudy (anonymous):

you have to find the value of each number and make them add up to the number at the end

OpenStudy (anonymous):

i understand what you have to do just busy right now and cant really help @SolomonZelman @ash2326 can yall help

OpenStudy (anonymous):

ok @KlOwNlOvE

OpenStudy (anonymous):

so @SolomonZelman and @ash2326 can y'all help? plz??

OpenStudy (anonymous):

@kropot72 ?

OpenStudy (anonymous):

@texaschic101 can you help?

OpenStudy (texaschic101):

do you know if there are any negative numbers used... the reason I ask, is if a + b + c = 16....look at the top row going across.. BCAC = 54....if a + b + c = 16, then the other c is 54 - 16 = 38...yet the total going down is 34. are the numbers supposed to toal going across, or just down ?

OpenStudy (texaschic101):

this is not making sense...the total of the numbers going across should equal the total of the numbers going down

OpenStudy (anonymous):

it confused me when i looked at it

OpenStudy (anonymous):

@Kimone are you there

OpenStudy (texaschic101):

well...there is 3 unknowns....so lets set up 3 equations A + B + 2C = 54 2A + B + C = 50 A + 2B + C = 44 multiply 1st eq by -2 : -2A - 2B - 4C = - 108 (result of multiplying by -2) 2A + B + C = 50 ---------------add -B - 3C = - 58 2A + B + C = 50 A + 2B + C = 44 --->(-2) -------------- 2A + B + C = 50 -2A - 4B - 2C = -88 (result of multiplying by -2) --------------add -3B - C = - 38 -B - 3C = -58 -3B - C = -38 --->(-3) ----------- -B - 3C = -58 9B + 3C = 114 -------------add 8B = 56 B = 56/8 B = 7 -3B - C = -38 -3(7) - C = -38 -21 - C = -38 -C = -38 + 21 -C = - 17 C = 17 A + B + 2C = 54 A + 7 + 2(17) = 54 A + 7 + 34 = 54 A + 41 = 54 A = 54 - 41 A = 13 B+ C + A + C = 54 7 + 17 + 13 + 17 = 54 54 = 54 A + C + B + A = 50 13 + 17 + 7 + 13 = 50 50 = 50 ok...I am confused now....I am getting them to equal going across...but I am not getting (A + B + C = 16) sorry...I don't know

OpenStudy (anonymous):

well thanks for trying and i was doing thesame thing until i got confused and stopped

OpenStudy (anonymous):

I'm here. I wen to the tutoring lad=b and the girl there said she had to stop too

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

This is killing me

OpenStudy (anonymous):

lol sorry this is confusing they didnt give any numbers?

OpenStudy (anonymous):

the only numbers are the ones at the end of each box

OpenStudy (anonymous):

hmm only thing i can say is talk to one of the moderators they may know ganeshi probably knows how

OpenStudy (anonymous):

@jdoe0001 ?

OpenStudy (anonymous):

Ok i'll definitely post the answer

OpenStudy (anonymous):

@phi can you try this?

OpenStudy (anonymous):

@ganeshie8 i know this was yesterday but can you try this question? He still hasn't been given help and i'm curious to see how it's solves